If 4.96 g Fe(NO3)3 is dissolved in enough water to make exactly 478 mL of solution, what is the molar concentration of nitrate ion?

A. 4.29×10-2 M
B. 1.43×10-2 M
C. 0.129 M
D. 306 M
E. 1.29×10-4 M
1
Expert's answer
2019-03-14T04:00:30-0400

Fe(NO3)3->Fe3++3NO3-

n(Fe(NO3)3)=4.96g/241.88g/mol=0.0205 mol

According to equation 1 mole if Fe(NO3)3 gives 3moles of NO3-, then n(NO3-)=3*n(Fe(NO3)3)= 0.0205*3=0.0615 mol

c(NO3-)= 0.0615mol/0.478L= 0.129 mol/L=0.129M

Answer:C.0.129M